Grade 11Physical Chemistry

calculate amount of CO2 that could be produced when 1). 1 mole of carbon is burnt in air 2). 1 mole of carbon is burnt in 16 gram of O2 3). 2 mole of Carbon is burnt in 16 gram of O2

Samay Dixit
7 Years agoGrade 11

2 Answers

Arun
7 Years ago

i) when 1 mole of carbon is burnt in 1 mole(32 g) of air(dioxygen) ,then according to chemical equation (1) 1mole(44 g) of carbon dioxide is produced.

 

ii) 32g of Oxygen upon heating gives CO2 = 44g

1 g of Oxygen upon heating gives CO2 = 44/32

 

Therefore 16 g of Oxygen upon heating gives CO2 = 44/32*16 = 22 g

 

(iii) According to the question, only 16 g of dioxygen is available. It is a limiting reactant. Thus, 16 g of dioxygen can combine with only 0.5 mole of carbon to give 22 g of carbon dioxide.

Manda Snehashish Reddy
7 Years ago
We know the chemical equation C + O2 = CO2. It is balanced.
  1. It can be seen from the equation that 1 mole of carbon gives rise to 1 mole of CO2 which is 44g (or) 22.4 Litres
  2. 16g of O2 is 0.5 mole. O2 acts as the limiting reagant here. Thus we obtain, 0.5 mole of CO2 only which is 22g (or) 11.2 Litres
  3. Still, O2 is the limiting reagant here. The obtained CO2 will be 22g (or) 11.2 Litres